# Corks, exotic 4-manifolds and genus functions

### v3

@@ 4. Producing infinite order generalized corks
The first example of an infinite order generalized cork was given by Tange [63, Theorem 1.6], and his example is simply connected and satisfies $b_{2}=2$. Gompf [36] later gave the first examples of infinite order contractible corks. We show that every 2-handlebody can be modified into an infinite order generalized cork, showing that infinite order generalized corks exist in abundance.
@@
For any $4$-dimensional $2$-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ topologically resembles $X$. Furthermore, in the case where $H_{2}(X)$ is a non-trivial group, there exist arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that $X_{i}$’s are pairwise exotic and topologically resemble $X$.
@@
We fix positive integers $r,s,m$ and abbreviate the contractible 4-manifold $C(r,s;m)$ in Subsection 3.5 by $C$, and we set $N=r+s+m-3$. Since $C$ is of Mazur type, its double $C\cup\overline{C}$ is diffeomorphic to $S^{4}$.
Let $H$ be a homology cobordism obtained from $I\times\partial X$ by taking a boundary sum with $C$, where the sum is taken outside the tubular neighborhood of the torus on which Gompf’s torus twist $\tau$ is applied. We easily see that the double $H\cup\overline{H}$ is diffeomorphic to $(I\times\partial X)\#(C\cup\overline{C})$ and hence to $I\times\partial X$. Thus, $H$ is an invertible homology cobordism from $\partial X$ to $\partial X\#\partial C$.
Let $X^{\prime}$ be the 4-manifold obtained from $X$ by attaching the invertible homology cobordism $H$. We easily see that $X^{\prime}$ deformation retracts to $X$ and is diffeomorphic to $X\natural C$. Let $f$ be the self-diffeomorphism of $\partial X^{\prime}$ that extends the torus twist $\tau$ of $\partial C$ as the identity on the rest of $\partial X^{\prime}$. Since $\tau$ extends to a self-homeomorphism of $C$, the self-diffeomorphism $f$ extends to a self-homeomorphism of $X^{\prime}$ that is identity on $X$. We thus see that this self-homeomorphism induces the identity map on $H_{2}(X^{\prime})$.
We will show that $(X^{\prime},f)$ is an infinite order generalized cork. By the assumption, $X$ admits an embedding into a symplectic 4-manifold $Y$ with weakly convex boundary. Let $E^{\prime}$ be the 4-manifold obtained from $E$ by attaching a 2-handle along a meridian of the 0-framed knot with sufficiently large negative framing. Then $Y\natural E^{\prime}$ is obtained from $Y$ by attaching 2-handles along a Legendrian link with the contact $(-1)$-framings, and thus $Y\natural E^{\prime}$ and its blow-ups $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ are symplectic 4-manifolds with weakly convex boundary ([25, Theorem 2.5], [66], see also [24, Theorem 5.8]). We here construct an embedding of $X^{\prime}$ into $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ by taking a boundary sum of $X\subset Y$ and $C\subset E^{\prime}$ using a thickened path from $\partial X$ to $\partial C$.
@@
For Proposition 4.2, we can relax the assumption that $X$ admits an embedding into symplectic 4-manifold with weakly convex boundary. Indeed, as seen from the proof, if $X\natural E$ admits an embedding into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ and non-vanishing Seiberg–Witten invariant such that the torus $T$ in $E$ represents an infinite order second homology class of $Z$, then the conclusion still holds.
@@ 5. Isomorphisms between symmetric bilinear forms
Throughout this section, we use the following setting. For each $i=1,2$, let $A_{i},B_{i}$ be finitely generated $\mathbb{Z}$-modules, and set $A^{\prime}_{i}=A_{i}\oplus B_{i}$. Let $p_{A_{i}}:A^{\prime}_{i}\to A_{i}$ and $p_{B_{i}}:A^{\prime}_{i}\to B_{i}$ denote the projections. Take an integral symmetric bilinear form $Q_{A_{i}}:A_{i}\times A_{i}\to\mathbb{Z}$, and define symmetric bilinear forms $Q_{B_{i}}:B_{i}\times B_{i}\to\mathbb{Z}$ and $Q_{i}:A^{\prime}_{i}\times A^{\prime}_{i}\to\mathbb{Z}$ by $Q_{B_{i}}(b,b^{\prime})=0$ and $Q_{i}(a+b,a^{\prime}+b^{\prime})=Q_{A_{i}}(a,a^{\prime})$, where $a,a^{\prime}\in A_{i}$ and $b,b^{\prime}\in B_{i}$. We assume that the symmetric bilinear form $Q_{\widetilde{A}_{i}}:(A_{i}/\mathrm{Tor})\times(A_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by each $Q_{A_{i}}$ is non-degenerate. These maps are models of intersection forms of 4-manifolds.
@@
Due to the definition of $Q_{i}$, we can check that $\mathrm{rank}\,A_{i}$ is equal to the maximal rank of a submodule $C_{i}$ of $A^{\prime}_{i}$ such that the symmetric bilinear form $(C_{i}/\mathrm{Tor})\times(C_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by $Q_{i}$ is non-degenerate. This fact implies $\mathrm{rank}\,A_{1}=\mathrm{rank}\,A_{2}$ and also $\mathrm{rank}\,B_{1}=\mathrm{rank}\,B_{2}$, since the bilinear forms on $A^{\prime}_{1}/\mathrm{Tor}$ and $A^{\prime}_{2}/\mathrm{Tor}$ are isomorphic. Due to the assumption that $A_{1},A_{2}$ or $B_{1},B_{2}$ are torsion-free, it follows that $A_{1},B_{1}$ are respectively isomorphic to $A_{2},B_{2}$.
Let each $T_{i}$ be the torsion submodule of $A_{i}$. Clearly, $T_{1}=T_{2}=0$ in the case where $A_{1},A_{2}$ are torsion-free. Let each $F_{i}$ be a free submodule of $A_{i}$ such that $A_{i}=F_{i}\oplus T_{i}$. We note that $F_{1},T_{1}$ are isomorphic to $F_{2},T_{2}$, respectively. Let $M_{i}$ be a matrix representing the bilinear form $F_{i}\times F_{i}\to\mathbb{Z}$ induced by $Q_{A_{i}}$. Then each $\det M_{i}$ is independent of the choices of $F_{i}$ and its basis, and $\det M_{i}$ is non-zero due to the non-degeneracy of $Q_{A_{i}}$ on $A_{i}/\mathrm{Tor}$.
Let each $p_{F_{i}}:A^{\prime}_{i}=F_{i}\oplus T_{i}\oplus B_{i}\to F_{i}$ be the projection. We here show that $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ is an isomorphism. Let $u_{1},\dots,u_{k}$ be a basis of $F_{1}$, and define $u_{1}^{\prime},\dots,u_{k}^{\prime}\in F_{2}$ by $u_{j}^{\prime}=p_{F_{2}}(\Phi(u_{j}))$. We define the matrix $M_{1}$ in the last paragraph by $M_{1}=(Q_{A_{1}}(u_{i},u_{j}))$. Since the projection $p_{F_{2}}$ preserves $Q_{2},Q_{A_{2}}$, the homomorphism $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ preserves $Q_{A_{1}},Q_{A_{2}}$, showing that the matrix $(Q_{2}(u_{i}^{\prime},u_{j}^{\prime}))$ is equal to $M_{1}$. Due to the condition $\det M_{1}\neq 0$, we easily see that $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is linearly independent (cf. [38, Corollary 1.2.13.]). We thus have a basis $v_{1},\dots,v_{k}$ of $F_{2}$ and positive integers $a_{1},\dots,a_{k}$ such that $a_{1}v_{1},\dots,a_{k}v_{k}$ generates the submodule generated by $u_{1}^{\prime},\dots,u_{k}^{\prime}$. It follows that $\det M_{1}=m\cdot\det M_{2}$ for some positive integer $m$. By reversing the roles of $A_{1}$ and $A_{2}$ and applying the same argument, we see that $\det M_{1}=\det M_{2}$. Consequently, $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is a basis of $F_{2}$, and thus $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ is an isomorphism.
@@
Finally, we show that $p_{B_{2}}\circ\Phi|_{B_{1}}:B_{1}\to B_{2}$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$. Due to the assumption that $A_{1},A_{2}$, or $B_{1},B_{2}$, are torsion-free, $\Phi$ bijectively maps the torsion submodule of $B_{1}$ to that of $B_{2}$. We can thus easily check $p_{B_{2}}\circ\Phi|_{B_{1}}(B_{1})$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$, by using the non-degeneracy assumption of each $Q_{A_{i}}$ on $A_{i}/\mathrm{Tor}$. ∎
In the rest of this section, let $R$ be an ordered set, and let $G_{i}:A^{\prime}_{i}\to R$ be an arbitrary map for each $i=1,2$. These maps are models of genus functions and related invariants of 4-manifolds. We will construct isomorphisms that preserve the restrictions of $Q_{1},Q_{2}$ and $G_{1},G_{2}$ by using Lemma 5.1.
@@
By Lemma 5.1, we see that $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ is an isomorphism that preserves $Q_{A_{1}}$, $Q_{A_{2}}$. Let $a$ be an element of $A_{1}$ and set $\widehat{a}=p_{A_{2}}\circ\Phi(a)$. We will prove $G_{1}(a)=G_{2}(\widehat{a})$. We see that $\Phi(a)=\widehat{a}+\widehat{a}_{B}$ for some $\widehat{a}_{B}\in B_{2}$. We consider two cases.
@@
Since $t$ is a torsion, this implies $G_{1}(a)=G_{2}(\widehat{a})$.
Therefore, the isomorphism $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ preserves $Q_{A_{1}}$, $Q_{A_{2}}$ and $G_{A_{1}},G_{A_{2}}$. ∎
We also discuss maps which are models of the torsion-free versions of the genus functions and related invariants. For each $i=1,2$, we set $\widetilde{A}_{i}=A_{i}/\mathrm{Tor}$, $\widetilde{B}_{i}=B_{i}/\mathrm{Tor}$ and $\widetilde{A}^{\prime}_{i}=\widetilde{A}_{i}\oplus\widetilde{B}_{i}$. Let $\widetilde{G}_{i}^{A}:\widetilde{A}_{i}\to R$, $\widetilde{G}_{i}^{B}:\widetilde{B}_{i}\to R$ and $\widetilde{G}_{i}:\widetilde{A}^{\prime}_{i}\to R$ be the maps defined by
$\widetilde{G}_{i}^{A}(\alpha)$ $=\min\{G_{i}(a)\mid a\in A_{i},\;[a]=\alpha\},$
$\widetilde{G}_{i}^{B}(\beta)$ $=\min\{G_{i}(b)\mid b\in B_{i},\;[b]=\beta\},$
$\widetilde{G}_{i}(\gamma)$ $=\min\{G_{i}(r)\mid r\in A_{i}^{\prime},\;[r]=\gamma\},$
where $[a]$, $[b]$ and $[r]$ denote their equivalent classes. We prove the following proposition.
Proposition 5.4.
Assume that each $G_{i}$ satisfies $G_{i}(a)\leq G_{i}(a+b)$ for any $a\in A_{i}$ and $b\in B_{i}$. If there exists an isomorphism $\Psi:\widetilde{A}^{\prime}_{1}\to\widetilde{A}^{\prime}_{2}$ that preserves $\widetilde{Q}_{1},\widetilde{Q}_{2}$ and $\widetilde{G}_{1},\widetilde{G}_{2}$, then there exists an isomorphism between $\widetilde{A}_{1}$ and $\widetilde{A}_{2}$ that preserves $Q_{\widetilde{A}_{1}},Q_{\widetilde{A}_{2}}$ and $\widetilde{G}_{1}^{A},\widetilde{G}_{2}^{A}$.
Proof.
Let $p_{\widetilde{A}_{i}}:\widetilde{A}^{\prime}_{i}\to\widetilde{A}_{i}$ and $p_{\widetilde{B}_{i}}:\widetilde{A}^{\prime}_{i}\to\widetilde{B}_{i}$ be the projections. By Lemma 5.1, the composition $p_{\widetilde{A}_{2}}\circ\Psi|_{\widetilde{A}_{1}}:\widetilde{A}_{1}\to\widetilde{A}_{2}$ is an isomorphism that preserves $Q_{\widetilde{A}_{1}}$ and $Q_{\widetilde{A}_{2}}$. For each $i=1,2$, we can easily check that every $\alpha\in\widetilde{A}_{i}$ and $\beta\in\widetilde{B}_{i}$ satisfy the inequality $\widetilde{G}_{i}(\alpha+\beta)\geq\widetilde{G}_{i}(\alpha)$ due to the assumption on $G_{i}$. This inequality implies $\widetilde{G}_{i}|_{\widetilde{A}_{i}}=\widetilde{G}^{A}_{i}$. Therefore, Proposition 5.3 shows that the above isomorphism preserves $\widetilde{G}_{1}^{A}$ and $\widetilde{G}_{2}^{A}$. ∎
@@ 6. Quasi-invertible cobordisms
In this section, we introduce the notion of quasi-invertible cobordisms to study genus functions. We also provide their examples and study their algebraic properties. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected. For such manifolds, the uniqueness of connected sums and boundary sums do not necessarily hold, but our results hold for any choices of sums unless otherwise stated. So, we will use ##\## and $\natural$ for these operations. Let $M,N$ be closed 3-manifolds.

### v5

@@ 4. Producing infinite order generalized corks
The first example of an infinite order generalized cork was given by Tange [65, Theorem 1.6], and his example is simply connected and satisfies $b_{2}=2$. Gompf [37] later gave the first examples of infinite order contractible corks. We show that every 2-handlebody can be modified into an infinite order generalized cork, showing that infinite order generalized corks exist in abundance.
@@
For any $4$-dimensional 22-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ is HIHC-equivalent to $X$. Furthermore, in the case where $H_{2}(X)\neq 0$, there exist arbitrarily many generalized corks of infinite order such that the $4$-manifolds constituting these corks are pairwise exotic and are HIHC-equivalent to $X$.
@@
We fix positive integers $r,s,m$ and abbreviate the contractible 4-manifold $C(r,s;m)$ in Subsection 3.5 by $C$, and we set N=r+s+m3N=r+s+m-3. Since $C$ is of Mazur type, its double $C\cup\overline{C}$ is diffeomorphic to $S^{4}$.
Let $H$ be a homology cobordism obtained from $I\times\partial X$ by taking a boundary sum with $C$, where the sum is taken outside the tubular neighborhood of the torus on which Gompf’s torus twist $\tau$ is applied. We see that the double $H\cup\overline{H}$ is diffeomorphic to $(I\times\partial X)\#(C\cup\overline{C})$ and hence to $I\times\partial X$. Thus, $H$ is an invertible homology cobordism from $\partial X$ to $\partial X\#\partial C$.
Let $X^{\prime}$ be the 4-manifold obtained from $X$ by attaching the invertible homology cobordism $H$. Then, $X^{\prime}$ deformation retracts to $X$ and is diffeomorphic to $X\natural C$. Let $f$ be the self-diffeomorphism of $\partial X^{\prime}$ that extends the torus twist $\tau$ of $\partial C$ as the identity on the rest of $\partial X^{\prime}$. Since $\tau$ extends to a self-homeomorphism of $C$, the self-diffeomorphism $f$ extends to a self-homeomorphism of $X^{\prime}$ that is identity on $X$. We thus see that this self-homeomorphism induces the identity map on $H_{2}(X^{\prime})$.
We will show that $(X^{\prime},f)$ is an infinite order generalized cork. By the assumption, $X$ admits an embedding into a symplectic 4-manifold $Y$ with weakly convex boundary. Let $E^{\prime}$ be the 4-manifold obtained from $E$ by attaching a 2-handle along a meridian of the 0-framed knot with sufficiently large negative framing. Then $Y\natural E^{\prime}$ is obtained from $Y$ by attaching 2-handles along a Legendrian link with the contact (1)(-1)-framings, and thus $Y\natural E^{\prime}$ and its blow-ups $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ are symplectic 4-manifolds with weakly convex boundary ([26, Theorem 2.5], [68], see also [25, Theorem 5.8]). We embed $X^{\prime}$ into $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ by taking a boundary sum of $X\subset Y$ and $C\subset E^{\prime}$ using a thickened path from $\partial X$ to $\partial C$.
@@
For each $X$, the above proof gives infinitely many pairwise non-homeomorphic 4-manifolds having the properties of $X^{\prime}$ in Proposition 4.2 by varying $r,s,m$ in the proof. Also, we can relax the assumption of Proposition 4.2 that $X$ admits an embedding into a symplectic 4-manifold with weakly convex boundary. Indeed, as seen from the proof, if $X\natural E$ admits an embedding into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ and non-vanishing Seiberg–Witten invariant such that the torus $T$ in $E$ represents an infinite order second homology class of $Z$, then the conclusion still holds.
@@ 5. Isomorphisms between symmetric bilinear forms
Throughout this section, we use the following setting. For each $i=1,2$, let $A_{i},B_{i}$ be finitely generated $\mathbb{Z}$-modules, and set $A^{\prime}_{i}=A_{i}\oplus B_{i}$. Let $p_{A_{i}}:A^{\prime}_{i}\to A_{i}$ and $p_{B_{i}}:A^{\prime}_{i}\to B_{i}$ denote the projections. Take an integral symmetric bilinear form $Q_{A_{i}}:A_{i}\times A_{i}\to\mathbb{Z}$, and define symmetric bilinear forms $Q_{B_{i}}:B_{i}\times B_{i}\to\mathbb{Z}$ and $Q_{i}:A^{\prime}_{i}\times A^{\prime}_{i}\to\mathbb{Z}$ by $Q_{B_{i}}(b,b^{\prime})=0$ and $Q_{i}(a+b,a^{\prime}+b^{\prime})=Q_{A_{i}}(a,a^{\prime})$, where $a,a^{\prime}\in A_{i}$ and $b,b^{\prime}\in B_{i}$. We set $\widetilde{A}_{i}=A_{i}/\mathrm{Tor}$ and assume that the symmetric bilinear form $Q_{\widetilde{A}_{i}}:\widetilde{A}_{i}\times\widetilde{A}_{i}\to\mathbb{Z}$ induced by each $Q_{A_{i}}$ is non-degenerate. These maps are models of intersection forms of 4-manifolds.
@@
Due to the definition of $Q_{i}$, we can check that $\mathrm{rank}\,A_{i}$ is equal to the maximum rank of a submodule $C_{i}$ of $A^{\prime}_{i}$ such that the symmetric bilinear form $(C_{i}/\mathrm{Tor})\times(C_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by $Q_{i}$ is non-degenerate. This fact implies $\mathrm{rank}\,A_{1}=\mathrm{rank}\,A_{2}$ and also $\mathrm{rank}\,B_{1}=\mathrm{rank}\,B_{2}$, since the bilinear forms on $A^{\prime}_{1}/\mathrm{Tor}$ and $A^{\prime}_{2}/\mathrm{Tor}$ are isomorphic. Due to the assumption that $A_{1},A_{2}$ or $B_{1},B_{2}$ are torsion-free, it follows that $A_{1},B_{1}$ are respectively isomorphic to $A_{2},B_{2}$.
Let each $T_{i}$ be the torsion submodule of $A_{i}$. Note that $T_{1}=T_{2}=0$ in the case where $A_{1},A_{2}$ are torsion-free. Let each $F_{i}$ be a free submodule of $A_{i}$ such that $A_{i}=F_{i}\oplus T_{i}$. We note that $F_{1},T_{1}$ are isomorphic to $F_{2},T_{2}$, respectively. Let $M_{i}$ be a matrix representing the bilinear form $F_{i}\times F_{i}\to\mathbb{Z}$ induced by $Q_{A_{i}}$. Then each $\det M_{i}$ is independent of the choices of $F_{i}$ and its basis, and $\det M_{i}$ is non-zero due to the non-degeneracy of $Q_{\widetilde{A}_{i}}$.
Let each $p_{F_{i}}:A^{\prime}_{i}=F_{i}\oplus T_{i}\oplus B_{i}\to F_{i}$ be the projection. We here show that $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ is an isomorphism. Let $u_{1},\dots,u_{k}$ be a basis of $F_{1}$, and define $u_{1}^{\prime},\dots,u_{k}^{\prime}\in F_{2}$ by $u_{j}^{\prime}=p_{F_{2}}(\Phi(u_{j}))$. We define the matrix $M_{1}$ in the last paragraph by $M_{1}=(Q_{A_{1}}(u_{i},u_{j}))$. Since the projection $p_{F_{2}}$ preserves $Q_{2},Q_{A_{2}}$, the homomorphism $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ preserves $Q_{A_{1}},Q_{A_{2}}$, showing that the matrix $(Q_{2}(u_{i}^{\prime},u_{j}^{\prime}))$ is equal to $M_{1}$. Due to the condition $\det M_{1}\neq 0$, we easily see that $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is linearly independent (cf. [39, Corollary 1.2.13.]). We thus have a basis $v_{1},\dots,v_{k}$ of $F_{2}$ and positive integers $a_{1},\dots,a_{k}$ such that $a_{1}v_{1},\dots,a_{k}v_{k}$ generates the submodule generated by $u_{1}^{\prime},\dots,u_{k}^{\prime}$. It follows that $\det M_{1}=m\cdot\det M_{2}$ for some positive integer $m$. By reversing the roles of $A_{1}$ and $A_{2}$ and applying the same argument, we see that $\det M_{1}=\det M_{2}$. Consequently, $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is a basis of $F_{2}$, and thus $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ is an isomorphism.
@@
Finally, we show that $p_{B_{2}}\circ\Phi|_{B_{1}}:B_{1}\to B_{2}$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$. Due to the assumption that $A_{1},A_{2}$, or $B_{1},B_{2}$, are torsion-free, $\Phi$ bijectively maps the torsion submodule of $B_{1}$ to that of $B_{2}$. We can thus easily check $p_{B_{2}}\circ\Phi|_{B_{1}}(B_{1})$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$, by using the non-degeneracy assumption of each $Q_{\widetilde{A}_{i}}$. ∎
In the rest of this section, let $R$ be an ordered set, and let $G_{i}:A^{\prime}_{i}\to R$ be an arbitrary map for each $i=1,2$. These maps are not necessarily homomorphisms and are models of genus functions and related invariants. We will construct isomorphisms that preserve the restrictions of $Q_{1},Q_{2}$ and $G_{1},G_{2}$ by using Lemma 5.1.
@@
[…]
@@
[…]
@@ 6. Quasi-invertible cobordisms
In this section, we introduce the notion of quasi-invertible cobordisms to study genus functions. We also provide their examples and study their algebraic properties. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected. For such manifolds, the uniqueness of connected sums and boundary sums do not necessarily hold, but our results hold for any choices of sums unless otherwise stated. So, we will use #\# and $\natural$ for these operations. Let $M,N$ be closed 3-manifolds.